OpenStudy (anonymous):

Jules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately what maximum height does the soccer ball reach?

OpenStudy (anonymous):

By Newton's third law of motion, v^2=u^2+2as where, "v" denotes final velocity, "u" denotes initial velocity, "a" denotes acceleration, and "s" denotes the height attained. Here, "v"=0. Therefore, (-u)^2=2as Or, -25*-25=2*10*s Or, 625=20s Or, s=625/20=31.25 feet

OpenStudy (anonymous):

??

OpenStudy (anonymous):

equation is \[h(t)=25t-16t^2\]

OpenStudy (anonymous):

maximum is given by \[\frac{-b}{2a}\] and here \[a=-16,b=25\] get \[\frac{-b}{2a}=\frac{25}{32}\]

OpenStudy (anonymous):

there shouldnt be anything wrong with using v^2 -u^2 +2as

OpenStudy (anonymous):

then plug it in

OpenStudy (anonymous):

I hope my equation and calculation is correct.

OpenStudy (anonymous):

maybe it is let me check

OpenStudy (anonymous):

thank youuu guys :)

OpenStudy (anonymous):

aadarsh the velocity is feet per second and you used m/s for gravity thats the only problem I think

OpenStudy (anonymous):

m/s^2

OpenStudy (anonymous):

i get \frac{1775}{64}=27.73 \]

OpenStudy (anonymous):

\[\frac{1775}{64}=27.73\]

OpenStudy (anonymous):

ya, ya!! ur correct. Let me do the correction. Or, -25*-25=2*10*10/3*s Or, 625=200/3s( Since, a=10 m/s^2, now, 1m=100 cm, and 30 cm=1 foot, so, 1 m=10/3 feet) Or, 625*3=200s Or, 1875=200s Or, s=1875/200=9.375 feet

OpenStudy (anonymous):

I did it both ways and got 625/64=9.77 feet both times

OpenStudy (anonymous):

using 32 ft/s^2 for gravity